# SQ-GAN: Semantic Image Communications Using Masked Vector Quantization

### v1

@@ Abstract
This work introduces Semantically Masked VQ-GAN (SQ-GAN), a novel approach integrating generative models to optimize image compression for semantic/task-oriented communications. SQ-GAN employs off-the-shelf semantic semantic segmentation and a new specifically developed semantic-conditioned adaptive mask module (SAMM) to selectively encode semantically significant features of the images. SQ-GAN outperforms state-of-the-art image compression schemes such as JPEG2000 and BPG across multiple metrics, including perceptual quality and semantic segmentation accuracy on the post-decoding reconstructed image, at extreme low compression rates expressed in bits per pixel. The code is available at https://github.com/frapez1/SQ-GAN
@@ I Introduction
The recent surge of intelligent and interconnected devices is leading the transition to new communication paradigms. The upcoming sixth generation (6G) of wireless networks is poised to redefine the communication paradigm by integrating advanced Artificial Intelligence (AI) and networking [1]. One of the most promising advancements in this context is semantic communication, which shifts the focus from raw data transmission to task-oriented and context-aware knowledge representation. Instead of transmitting all information indiscriminately, semantic communication prioritizes the meaning and relevance of the data, for a given application, and enhancing communication efficiency [2], [3, 4, 5, 6].
Recent breakthroughs in deep learning, generative AI, and self-supervised learning have further strengthened the potential of semantic communications. The use of generative AI for next generation networks has been recently proposed in [7], [8], [9]. A key innovation in this domain is the integration of deep generative models that can synthesize high-fidelity multimedia content once properly conditioned. This capability shifts the communication paradigm: rather than transmitting full-resolution data, the transmitter selects and encodes only the most essential semantic features needed to trigger the generative model at the receiver side [10]. The receiver then reconstructs a meaningful representation, conditioned on the received information, and tailored to the application at hand.
@@
To illustrate this concept, consider a vehicular communication scenario, where an autonomous vehicle transmits visual information from its onboard cameras to a roadside unit (RSU) responsible for making real-time context-aware safety decisions, possibly integrating data from multiple vehicles or RSU sensors. Instead of sending full video streams, the transmitter extracts key semantic attributes, such as the locations of pedestrians, vehicles, traffic signs, and road boundaries, and transmits compressed feature representations. At the receiver side, a generative model reconstructs an application-specific representation, ensuring that the RSU can accurately interpret the environment and make fast, informed decisions.
@@
Deep learning has played a pivotal role in advancing semantic communication systems [11]. The development of end-to-end learning frameworks enables joint optimization of encoding and decoding processes, allowing systems to learn joint source/channel coding [12] or learn efficient representations of semantic content directly from data [13, 14]. Generative models, particularly Generative Adversarial Networkss (GANs), have been instrumental in improving the quality and efficiency of data representation and transmission [15, 16]. Specifically, integrating vector quantization techniques with generative models has shown great potential in learning discrete latent representations that are both compact and semantically rich [17, 18]. Models like the Vector Quantized VAE (VQ-VAE), the Vector Quantized GAN (VQ-GAN) and the Masked VQ-VAE (MQ-VAE) [19] have demonstrated remarkable capabilities in image synthesis and compression tasks.
In this context, we propose the Semantic Masked VQ-GAN (SQ-GAN) model, a novel VQ-GAN-based approach that merges the strengths of generative models and semantic communication for efficient data transmission. This architecture is designed to compress and reconstruct an image $\mathbf{x}$ such that its Semantic Segmentation Map (SSM) $\mathbf{s}$ is preserved. The SQ-GAN makes use of masked vector quantization [19] to prioritize semantically significant information, thereby reducing redundancy and enhancing communication efficiency.
The SQ-GAN framework operates by selectively encoding the most relevant features of the input data based on their semantic importance. This selective encoding is achieved through a newly designed and trained adaptive masking mechanism that learns to identify and preserve critical information while reducing redundancy. The masked vector quantization not only compresses data effectively but also ensures that semantically important information is preserved, making it highly suitable for applications where bandwidth and computational resources are limited. The overall scheme is a hierarchical coding scheme, where different objects in the image have different importance, depending on the task subsuming the exchange of information. The proposed approach can be seen as a generalization of hierarchical image coding [20], where the linear transform stage, e.g. a wavelet transform, is replaced by a non-linear learned “encoder” stage, guided by the segmentation map of the image, obtained after prior selection of the classes of objects deemed relevant for the application. The encoder produces a real-valued latent space tensor, which plays the role of the transform coefficients in conventional image coding schemes. The new adaptive masking mechanism is trained to select the elements in the latent space tensor that are most significant for the preservation of the image semantic content. Only these selected elements are then quantized, using vector quantization with the learned codebook. Finally, the decoder part of the scheme is formed by some learned blocks, trained to reconstruct the original image by minimizing an error function that includes its SSM.
@@
The key innovations in the proposed SQ-GAN scheme are: (i) a novel GAN-based model that integrates semantic conditioning and compression directly into its architecture, (ii) the Semantic conditioned Adaptive Mask Module (SAMM) that selects and encodes the most semantically relevant features of the input data based on the SSM, (iii) a specifically designed data augmentation technique in the training phase to enhance the semantically relevant classes, and (iv) the inclusion (in training) of a semantic-aware discriminator to force the model to give more importance to the semantically relevant regions over the non-relevant ones.
This work is organized as follows. Section II introduces the scope and general overview of the proposed architecture. Section III introduces in more detail the different parts of the proposed SQ-GAN model. Section IV discusses the training method with data augmentation for the enhancement of semantic relevant classes and the semantic-aware discriminator employed to optimize the SQ-GAN. Finally, Section V presents numerical experiments demonstrating the effectiveness of SQ-GAN compared to state-of-the-art image compression techniques such as (JPEG2000) and Better Portable Graphics (BPG).
II Problem definition and general architecture
The objective of this work is to develop an image encoder that efficiently compresses an image $\mathbf{x}$ along with its semantic segmentation map $\mathbf{s}$, in order to enable the reconstruction of an image $\hat{\mathbf{x}}$ that preserves the semantic content of the original image, even at very low Bits Per Pixel (BPP) (see Fig. 1). The model leverages both the original image $\mathbf{x}$ and its corresponding SSM $\mathbf{s}$, which provides pixel-wise class labels indicating the semantic class of each pixel in the image. The SSM $\mathbf{s}$ is obtained using a state-of-the-art (SOTA) Semantic Segmentation Model (SSModel) [21].
The encoder in SQ-GAN jointly processes $\mathbf{x}$ and $\mathbf{s}$ and produces two bit streams, which are then used to reconstruct $\hat{\mathbf{x}}$ and $\hat{\mathbf{s}}$. To evaluate the semantic quality of $\hat{\mathbf{x}}$, we also define the “generated” SSM $\tilde{\mathbf{s}}$, which is obtained by applying a a SOTA SSModel (e.g., the same which was used to generate $\mathbf{s}$ from $\mathbf{x}$) to the reconstructed image $\hat{\mathbf{x}}$. Notice that, in general, $\mathbf{s}$, $\hat{\mathbf{s}}$, and $\tilde{\mathbf{s}}$, are different.
By jointly encoding and decoding $\mathbf{x}$ and $\mathbf{s}$, and incorporating the reconstruction of $\hat{\mathbf{s}}$ in the decoder, it is possible to directly optimize the preservation of semantic information during the training process and use $\hat{\mathbf{s}}$ to condition the reconstruction $\hat{\mathbf{x}}$. This conditioning “pushes” the decoder to align $\hat{\mathbf{x}}$ with the semantic structure of the original image $\mathbf{x}$.
As anticipated above, the “generated” SSM $\tilde{\mathbf{s}}$ is obtained by applying a pre-trained SOTA SSModel to the reconstructed image $\hat{\mathbf{x}}$. By comparing $\tilde{\mathbf{s}}$ to the original $\mathbf{s}$, we can evaluate how well the reconstructed image $\hat{\mathbf{x}}$ retains semantic information recognizable by external models.
@@ II Problem definition and general architecture
The overall architecture of the proposed SQ-GAN scheme is shown in Fig. 2, where the original image $\mathbf{x}$ and the corresponding original SSM $\mathbf{s}$ are both represented by 3-dimensional tensors. The image $\mathbf{x}$ is a tensor of size $3\times H\times W$, where $3$ refers to the RGB channels, $H$ is the height and $W$ is the width of the frame (in pixels). In this work, we have used images of size $H=256$ and $W=512$. The SSM $\mathbf{s}$ is a tensor of size $n_{c}\times H\times W$, where $n_{c}$ refers to the number of semantic classes, $H$ is the height and $W$ is the width of the frame. Each $n_{c}$-dimensional vector in pixel position $(h,w)$ represents the 1-hot encoding of the corresponding semantic class of the pixel. This representation is quite standard in this setting and has proven to be convenient in a large number of works on semantic-conditioned image generation.
@@
The compression and reconstruction of the image $\mathbf{x}$ and of its SSM $\mathbf{s}$ is accomplished by two parallel (and mutually interacting) pipelines. We use apex/subscripts $\mathbf{x}$ and $\mathbf{s}$ to indicate blocks of the specific pipeline, in Fig. 2 they are represented by the beige and green color, respectivelly. In particular, the scheme is formed by two encoder-decoder pairs, $E_{\mathbf{x}}$-$D_{\mathbf{x}}$ and $E_{\mathbf{s}}$-$D_{\mathbf{s}}$, respectively. The semantic encoder $E_{\mathbf{s}}$ take as input the original SSM $\mathbf{s}$ and outputs a latent tensor $\mathbf{z}^{\mathbf{s}}=E_{\mathbf{s}}(\mathbf{s})$ of shape $C\times H_{16}\times W_{16}$ where $C=256$, $H_{16}=\frac{H}{16}$ and $W_{16}=\frac{W}{16}$. At the same time the SSM $\mathbf{s}$ and the image $\mathbf{x}$ are combined by the proposed Semantic Positional Embedding (SemPE) to influence the feature extraction of the image with the SSM. This new representation $\mathbf{h}_{SemPE}(\mathbf{x},\mathbf{s})$ is then encoded by $E_{\mathbf{x}}$ to obtain a latent tensor $\mathbf{z}^{\mathbf{x}}=E_{\mathbf{x}}(\mathbf{h}_{SemPE}(\mathbf{x},\mathbf{s}))$ with the same shape as $\mathbf{z}^{\mathbf{s}}$. Both tensors $\mathbf{z}^{\mathbf{x}}$ and $\mathbf{z}^{\mathbf{s}}$ are composed of $K=H_{16}\times W_{16}$ latent vectors $\mathbf{z}_{k}^{\mathbf{x}}$ and $\mathbf{z}_{k}^{\mathbf{s}}$ in $C$ dimensions.
@@
The SAMM blocks select the most semantically relevant latent vectors. The number of selected latent vectors depends on a masking fraction variable $m^{\mathbf{x}}$, or $m^{\mathbf{s}}$, that can be set as a parameter. Hence, only $N_{\mathbf{x}}=m^{\mathbf{x}}K$ and $N_{\mathbf{s}}=m^{\mathbf{s}}K$ selected latent vectors from $\mathbf{z}^{\mathbf{x}}$ and $\mathbf{z}^{\mathbf{s}}$ are quantized, respectively.
@@
Vector quantization (VQ) is applied separately to each selected vector, where the quantization codebook is formed by $J$ learnable $C$-dimensional vectors (quantization codewords) denoted by $\mathbf{e}_{j}^{\mathbf{x}}$ and $\mathbf{e}_{j}^{\mathbf{s}}$, $j\in\{1,\ldots,J\}$, respectively. Therefore, each quantized latent vector is identifies by a binary index of $\log_{2}J$ bits. In this work we chose $J=1024$ and the ($\ell_{2}$) distance as the quantization metric (more in Section III-D. The quantized latent tensors are denoted by $\mathbf{z}_{q}^{\mathbf{x}}$ and $\mathbf{z}_{q}^{\mathbf{s}}$, respectively. The list of quantization indices $\mathbf{e}_{j}^{\mathbf{x}}$ and $\mathbf{e}_{j}^{\mathbf{s}}$, including the positions of the discarded non-relevant indices, are binary encoded and from the binary encoded stream to be stored or transmitted.
@@
At the receiver side, the binary stream is decoded and formatted into latent space tensors $\mathbf{z}_{q}^{\mathbf{x}}$ and $\mathbf{z}_{q}^{\mathbf{s}}$. Following the same procedure as in [19], the latent tensors contain the quantized vectors in the positions where these vectors were selected and effectively quantized, and a “placeholder” codeword $M_{\mathbf{x}}$ or $M_{\mathbf{s}}$ in the positions these were discarded. These placeholder codewords are not necessarily all-zero vectors, and can be learned in the traning phase in order to faclitate reconstruction. These latent space tensors are then processed by the Adaptive De-Mask Module (ADM) module to produce the reconstructed latent space tensors $\hat{\mathbf{z}}^{\mathbf{x}}$ and $\hat{\mathbf{z}}^{\mathbf{s}}$, respectively.
@@
In the next sections we will examine the detailed architecture of the blocks in Fig. 2, and the overall training method. In particular, it is important to clarify that the training of the proposed SQ-GAN is performed via an adversarial approach based on two discriminators. However, in Fig. 2 only the structure of the generator is represented. In fact, the generator is the part used at the running time, while the discriminators will only play a role in training.

### v2

@@ Abstract
This work introduces Semantically Masked Vector Quantized Generative Adversarial Network (SQ-GAN), a novel approach integrating semantically driven image coding and vector quantization to optimize image compression for semantic/task-oriented communications. The method only acts on source coding and is fully compliant with legacy systems. The semantics is extracted from the image computing its semantic segmentation map using off-the-shelf software. A new specifically developed semantic-conditioned adaptive mask module (SAMM) selectively encodes semantically relevant features of the image. The relevance of the different semantic classes is task-specific, and it is incorporated in the training phase by introducing appropriate weights in the loss function. SQ-GAN outperforms state-of-the-art image compression schemes such as JPEG2000, BPG, and deep-learning based methods across multiple metrics, including perceptual quality and semantic segmentation accuracy on the reconstructed image, at extremely low compression rates. The code is available at https://github.com/frapez1/SQ-GAN.
@@ I Introduction
The recent surge of intelligent and interconnected devices is leading the transition to new communication paradigms. The upcoming sixth generation (6G) of wireless networks is poised to redefine the communication paradigm by integrating advanced Artificial Intelligence (AI) and networking [1]. One of the most promising advancements in this context is semantic communication, which shifts the focus from universal coding to task-oriented coding, exploiting suitable knowledge representation systems, typically relying on deep neural networks [2], [3, 4, 5, 6]. Recent breakthroughs in deep learning, generative AI, and self-supervised learning have further strengthened the potential of semantic communications. The use of generative AI for next generation networks has been recently proposed in [7], [8], [9]. A key innovation in this domain is the integration of deep generative models that can synthesize, with proper conditioning, high-fidelity multimedia contents. This capability changes the communication paradigm: instead of transmitting full-resolution data, the transmitter encodes and sends only the most relevant features, which are used at the receiver side to trigger the generative model to reproduce a representation semantically equivalent to the transmitted one [10]. The relevance of the features is dictated by the application running on top of the exchange of information. This communication modality enables a drastic reduction in transmission rates, because it takes advantage of application-specific knowledge incorporated in the generative models.
@@
To illustrate this concept, consider a vehicular communication scenario, where one or more autonomous vehicles transmit visual information from their onboard cameras to a roadside unit (RSU) responsible for taking real-time context-aware safety decisions, possibly integrating data from multiple vehicles or RSU sensors. Rather than transmitting full video streams, each transmitter extracts key semantic attributes, such as the presence, shape, and locations of pedestrians, vehicles, traffic signs, and road boundaries, and transmits only compressed feature representations, achieving data rates substantially lower than those needed for conventional image or video encoding. At the receiver side, a generative model reconstructs an application-specific representation, with an accuracy sufficient only to enable the RSU to take rapid decisions within strict latency constraints.
@@
In recent years, deep learning has played a pivotal role in both improving image compression and advancing semantic communication systems [11, 12, 13]. The development of end-to-end learning frameworks enables the joint optimization of the encoding and decoding processes, allowing the overall encoding system to learn a joint source/channel coding (JSCC) technique [14] or to learn efficient representations of semantic content directly from data [15, 11]. Lately, generative models, such as Variational-Autoencoders (VAE), Generative Adversarial Networks (GANs), and Denoising Diffusion Probabilistic Models (DDPM) [16], have been instrumental in improving the quality and efficiency of data representation and transmission [7, 10, 17, 18].
A key step in the integration of generative models within a digital communication scheme, possibly compliant with legacy systems, is to merge generative models and vector quantization to learn discrete latent representations that are both compact and semantically rich [19, 20]. Models like the Vector Quantized VAE (VQ-VAE), the Vector Quantized GAN (VQ-GAN) and the Masked VQ-VAE (MQ-VAE) [21] have already demonstrated remarkable capabilities in image synthesis and compression tasks.
Given this context, in this paper we propose the Semantic Masked VQ-GAN (SQ-GAN) model, a novel approach that merges the strengths of generative models, vector quantization, and semantic segmentation for efficient and semantic-preserving image transmission. We choose the GAN paradigm because it provides a good tradeoff between simplicity and accuracy. In the proposed scheme, semantic information is extracted from the image $\mathbf{x}$ to be transmitted by computing its Semantic Segmentation Map (SSM) $\mathbf{s}$, using some suitable state-of-the-art (SOTA) algorithm (in our implementation we used INTERN-2.5 [22]). Then, SQ-GAN applies a semantically-guided vector quantization strategy to encode both $\mathbf{x}$ and $\mathbf{s}$. In particular, SQ-GAN learns a vector quantization codebook from the data and applies a newly designed and trained adaptive selective masking to dynamically adjust the compression rate, ensuring that semantically important features are retained. This is obtained by a specifically designed semantic-conditioned variant of the masked vector quantization introduced in [21] to prioritize semantically significant information, thus reducing redundancy while preserving the relevant semantic content. The new adaptive masking mechanism is trained to select the elements in the latent space tensor that are most significant for the preservation of the image’s semantic content. Only these selected elements are then quantized, using vector quantization with the learned codebook. Finally, the decoder part of the scheme is formed by some learned blocks, trained to reconstruct the original image by minimizing an error function that includes the SSM. The key innovations in the proposed SQ-GAN scheme are: (i) a novel GAN-based model that integrates semantic conditioning and compression directly into its architecture, (ii) the Semantic conditioned Adaptive Mask Module (SAMM) that selects and encodes the most semantically relevant features of the input data based on the SSM, (iii) a specifically designed data augmentation technique in the training phase to enhance the semantically relevant classes, and (iv) the inclusion (in training) of a semantic-aware discriminator to force the model to give more importance to the semantically relevant regions over the non-relevant ones.
Before proceeding, some important remarks regarding our method are in order:
i) Since we focus on task-oriented semantic communication, we assume that the semantic classes used by the state-of-the-art semantic segmentation algorithm are predefined according to the task requirements. Furthermore, the relevance of each class is quantified by weights $w_{i}\in[0,1]$ assigned to each $i$-th class. These weights are incorporated into the loss functions used to train the proposed SQ-GAN. The combination of this task-oriented weighted loss function and the aforementioned task-specific data augmentation during training results in an SQ-GAN model optimized for the target application. For instance, in our numerical examples focused on autonomous driving scenarios, the relevant object classes include pedestrians, cars, traffic lights, and similar entities.
ii) The assumption that the SSM $\mathbf{s}$ is extracted from the raw image $\mathbf{x}$ in real-time at the encoder is fully realistic. For example, in the autonomous driving scenario, there are algorithms able to extract the SSM in real time, see, e.g. [23]. In particular, Tesla has already deployed the technology to analyze raw images and perform semantic segmentation and object detection in their per-camera networks in real time, as reported in [24].
iii) Our scheme treats only source coding (i.e., image compression/reconstruction), unlike several works on semantic communications that have considered a JSCC setting. This makes our approach completely compatible with any underlying legacy network protocol, where source coding/decoding takes place at the application layer and the source-coded bits are transmitted to the destination via some standard protocol stack (e.g., a 3GPP wireless network).
This work is organized as follows. Section II introduces the scope and general overview of the proposed architecture. Section III introduces in more detail the different parts of the proposed SQ-GAN model. Section IV discusses the training method with data augmentation for the enhancement of semantically relevant classes and the semantic-aware discriminator employed to optimize the SQ-GAN. Finally, Section V presents numerical experiments demonstrating the effectiveness of SQ-GAN compared to state-of-the-art (SOTA) image compression techniques such as JPEG2000, Better Portable Graphics (BPG) and the GAN based High-Fidelity Generative Image Compression (HiFiC) [12] and Fidelity-Controllable extreme image Compression (FCC) [13].
II Problem statement and general architecture
@@
The objective of this work is to develop an image encoder that efficiently compresses an image $\mathbf{x}$, while preserving its semantic content as much as possible, even at extremely high compression rates. In this work, the semantic content is identified with the SSM $\mathbf{s}$, extracted in real-time from $\mathbf{x}$ using a SOTA Semantic Segmentation Model (SSModel) [22] with pre-defined semantic classes that depend on the task at hand, as explained in Section I.
As schematically shown in Fig. 1, our encoding scheme extracts $\mathbf{s}$ from $\mathbf{x}$ and then jointly encodes $\mathbf{x}$ and $\mathbf{s}$. At the decoder, $\mathbf{x}$ and $\mathbf{s}$ are reconstructed as $\hat{\mathbf{x}}$ and $\hat{\mathbf{s}}$, respectively. In some sense, this approach is akin layered source coding, where $\mathbf{s}$ can be seen as the “fundamental layer” and $\mathbf{x}$ is the refinement. The goal is to achieve good reconstruction performance (with respect to distortion metrics to be defined later) at a low overall coding rate, in terms of Bits Per Pixel (BPP). The key result of our work is to show that, even though we transmit two images, namely $\mathbf{x}$ and $\mathbf{x}$ instead of just $\mathbf{x}$, we achieve a significantly better trade-off between the overall coding rate and various metric for semantic distortion with respect to SOTA schemes (including data-driven approaches).
By jointly encoding and decoding $\mathbf{x}$ and $\mathbf{s}$, and incorporating the reconstruction of $\hat{\mathbf{s}}$ in the decoder, it is possible to directly optimize the preservation of semantic information and assign different importance to different semantic classes during the training process and use $\hat{\mathbf{s}}$ to condition the reconstruction of $\hat{\mathbf{x}}$. Intuitively, this conditioning “pushes” the decoder to align $\hat{\mathbf{x}}$ with the semantic structure of the original image $\mathbf{x}$.
@@
The detailed architecture of the proposed SQ-GAN scheme is shown in Fig. 2, where the original image $\mathbf{x}$ and the corresponding original SSM $\mathbf{s}$ are both represented by 3-dimensional tensors. The image $\mathbf{x}$ is a tensor of size $3\times H\times W$, where $3$ refers to the RGB channels, $H$ is the height and $W$ is the width of the frame (in pixels). In this work, we have used images of size $H=256$ and $W=512$. The SSM $\mathbf{s}$ is a tensor of size $n_{c}\times H\times W$, where $n_{c}$ is the number of semantic classes, $H$ is the height and $W$ is the width of the frame. Each $n_{c}$-dimensional vector in pixel position $(h,w)$ represents the 1-hot encoding of the corresponding semantic class of that pixel. This representation is quite standard and has proven to be convenient in a large number of works on semantic-conditioned image generation.
@@
The compression and reconstruction of the image $\mathbf{x}$ and of its SSM $\mathbf{s}$ is accomplished by two parallel (and mutually interacting) pipelines. We use apex/subscripts $\mathbf{x}$ and $\mathbf{s}$ to indicate blocks of the specific pipeline. Looking at the transmitter side, the bottom pipeline (in beige) reports the encoder of the SSM, whereas the top pipeline (in green) reports the image encoding process, which is conditioned to the SSM. The semantic encoder $E_{\mathbf{s}}$ takes as input the original SSM $\mathbf{s}$ and outputs the latent tensor $\mathbf{z}^{\mathbf{s}}=E_{\mathbf{s}}(\mathbf{s})$ of shape $C\times H_{16}\times W_{16}$ where $C=256$, $H_{16}=\frac{H}{16}$ and $W_{16}=\frac{W}{16}$. In parallel, the image encoding pipeline first extracts the latent representations of $\mathbf{x}$ and $\mathbf{s}$, using two pre-trained DNNs, namely the ${\rm InputNet}_{\mathbf{x}}$ for $\mathbf{x}$ and the proposed Semantic Positional Embedding (SemPE) for $\mathbf{s}$. These latent representations are added together as $\mathbf{h}(\mathbf{x},\mathbf{s})={\rm InputNet}_{\mathbf{x}}(\mathbf{x})+{\rm SemPE}(\mathbf{s})$ and sent to the encoder that produces the latent tensor $\mathbf{z}^{\mathbf{x}}=E_{\mathbf{x}}(\mathbf{h}(\mathbf{x},\mathbf{s}))$, having the same shape as $\mathbf{z}^{\mathbf{s}}$. Both tensors $\mathbf{z}^{\mathbf{x}}$ and $\mathbf{z}^{\mathbf{s}}$ are composed of $K=H_{16}\times W_{16}$ latent vectors $\mathbf{z}_{k}^{\mathbf{x}}$ and $\mathbf{z}_{k}^{\mathbf{s}}$ in $C$ dimensions.
@@
The tensors $\mathbf{z}^{\mathbf{x}}$ and $\mathbf{z}^{\mathbf{s}}$ are then sent to the corresponding Semantic conditioned Adaptive Mask Module (SAMM) blocks, which select the most semantically relevant latent vectors. The SAMM blocks play a key role in reducing the transmission rate, while preserving the most relevant information. The number of selected latent vectors depends on a masking fraction variable $m^{\mathbf{x}}$, or $m^{\mathbf{s}}$, that can be set as a parameter. Hence, only $N_{\mathbf{x}}=\left\lfloor m^{\mathbf{x}}K\right\rfloor$ and $N_{\mathbf{s}}=\left\lfloor m^{\mathbf{s}}K\right\rfloor$ selected latent vectors from $\mathbf{z}^{\mathbf{x}}$ and $\mathbf{z}^{\mathbf{s}}$ are quantized, respectively.
@@
Vector quantization (VQ) is applied separately to each selected vector, where the quantization codebook is formed by $J$ learnable $C$-dimensional vectors (quantization codewords) denoted by $\mathbf{e}_{j}^{\mathbf{x}}$ and $\mathbf{e}_{j}^{\mathbf{s}}$, $j\in\{1,\ldots,J\}$, respectively. Therefore, each quantized latent vector is identified by a binary index of $\log_{2}J$ bits. In this work we chose $J=1024$ and the $\ell_{2}$ distance as the quantization metric (more in Section III-D). The quantized latent tensors are denoted by $\mathbf{z}_{q}^{\mathbf{x}}$ and $\mathbf{z}_{q}^{\mathbf{s}}$, respectively. The list of quantization indices $\mathbf{e}_{j}^{\mathbf{x}}$ and $\mathbf{e}_{j}^{\mathbf{s}}$, including the positions of the discarded non-relevant indices, are binary encoded and form the binary encoded stream to be transmitted.
@@
At the receiver side, the binary stream is decoded and formatted into the latent space tensors $\mathbf{z}_{q}^{\mathbf{x}}$ and $\mathbf{z}_{q}^{\mathbf{s}}$. Following a similar procedure as in [21], the latent tensors contain the quantized vectors in the positions where these vectors were selected and effectively quantized, and a “placeholder” codeword $M_{\mathbf{x}}$ or $M_{\mathbf{s}}$ in the masked-out positions. These placeholder codewords are not necessarily all-zero vectors, and can be learned in the training phase in order to facilitate reconstruction. These latent space tensors are then processed by the Adaptive De-Mask Module (ADM) module to produce the reconstructed latent space tensors $\hat{\mathbf{z}}^{\mathbf{x}}$ and $\hat{\mathbf{z}}^{\mathbf{s}}$, respectively.
@@
The details of the blocks appearing in Fig. 2 and the overall training method will be illustrated in the next section. In particular, it is important to clarify that the training of the proposed SQ-GAN is performed via an adversarial approach based on two discriminators. However, in Fig. 2 only the structure of the generator is represented. In fact, the generator is eventually the trained model deployed and used at the run-time, while the discriminators play a role only in the training phase.